OpenStudy (anonymous):

four more than twice a number is 8

OpenStudy (tkhunny):

Please give the mystery number a name. Anything. Just name it.

OpenStudy (anonymous):

4+2x=8

OpenStudy (nfcfox):

A number you can call a variable, such as x. @plzhelpmethankyou Created an equation based on this.

OpenStudy (nfcfox):

And to solve it let's solver for x, our number. Subtract 4 so we get 2x=8-4 which is 2x=4. Now we divide the 4 by 2. x=4/2 which is x=2. Our number is 2!

OpenStudy (tkhunny):

YOu can call it "Steve" if you like. Just give it a name!